It IS BELIEVED THAT 1/4 OF ADULTS OVER 50 YEARS OLD NEVER GRADUATE FROM HIGH SCHOOL. We  WANT TO SEE IF THE PERCENTAGE IS THE SAME FOR 30 YEAR olds. How MANY Should WE SURVEY IN ORDER TO ESTIMATE THE PROPORTION OF NON-GRADS WITHIN 0.06 WITH 90% CONFIDENCE?
Answer #1

given data,
sample proportion is 50%
margin of error = 0.06
confidence level is 90%
sample size is
Compute Sample Size ( n ) = n=(Z/E)^2*p*(1-p)
Z a/2 at 0.1 is = 1.645
Sample Proportion = 0.25
ME = 0.06
n = ( 1.645 / 0.06 )^2 * 0.25*0.75
= 140.9388 ~ 141